# SparseInfer: Training-free Prediction of Activation Sparsity for Fast LLM Inference

### v1

@@ Abstract
Leveraging sparsity is crucial for optimizing large language model (LLM) inference; however, modern LLMs employing SiLU as their activation function exhibit minimal activation sparsity.
Recent research has proposed replacing SiLU with ReLU to induce significant activation sparsity and showed no downstream task accuracy degradation through fine-tuning.
However, taking full advantage of it required training a predictor to estimate this sparsity. In this paper, we introduce SparseInfer, a simple, light-weight, and training-free predictor for activation sparsity of ReLU-fied LLMs, in which activation sparsity is predicted by comparing only the sign bits of inputs and weights. To compensate for possible prediction inaccuracy, an adaptive tuning of the predictor’s conservativeness is enabled, which can also serve as a control knob for optimizing LLM inference. The proposed method achieves approximately faster inference speed over the state-of-the-art, with negligible accuracy loss of within 1%p.
@@ I Introduction
Following the remarkable success of Generative Pre-trained Transformer (GPT) models [1] based on the Transformer architecture [2], Large Language Models (LLMs) are being actively employed across a wide range of fields. The inference of LLMs based on the GPT architecture first performs an initial forward pass, also known as prefill, followed by an autoregressive decoding process that generates tokens sequentially, one at a time. Due to this sequential nature, memory accesses occur frequently in the decoding stage. This can lead to I/O bottlenecks, which in turn limits the overall computational performance. Well-known optimization methods such as pruning [3] or quantization [4] have been applied to LLMs to reduce this memory access overhead, but these approaches inevitably result in a loss of inference accuracy.
@@
The contributions of this paper are summarized as follows:
  • Unlike existing methods, the proposed technique leverages activation sparsity without requiring additional training.
  • The proposed method is efficiently implemented for GPU execution, and further optimized through kernel fusion across multiple stages of the MLP block, demonstrating an improvement in inference speed over the state-of-the-art, while maintaining minimal accuracy degradation.
  • The proposed predictor is designed to be tunable, allowing per-layer adjustment of the conservativeness of prediction to balance the trade-off between speed and accuracy.
@@ II Related Work
Activation Sparsity-induced LLM: Mirzadeh et al.[7] proposed a technique called ReLUfication which replaces activation functions in LLM, such as SiLU, with ReLU. They showed ReLUfication can induce significant activation sparsity while accuracy loss is negligible after fine-tuning. ProSparse [8] significantly induced activation sparsity by progressively applying sparsity regularization after Relufication. It also set a positive value instead of zero for a ReLU threshold and achieves larger sparsity using a technique called FATReLU [11]. It induced a 22%p higher sparsity compared to ReluLLaMa [12], an open-source ReLU-based LLM fine-tuned from Llama2, achieving 66% sparsity. Despite a slight performance drop in some common sense benchmarks, the resultant model generally maintained comparable performance on average and even performed slightly better in certain cases, achieving up to a 4.65×4.65$\times$4.65 × speedup in inference time with PowerInfer [13]. A similar approach, TurboSparse [14], was proposed for Mixtral models.
@@ III Background: Gate-Based MLP
The gate-base MLP block consists of three linear layers ($W_{gate}$, $W_{up}$, and $W_{down}\in\mathbb{R}^{d\times k}$ in which $k>d$) with $d$ being the model dimension and one activation layer ($\sigma(\cdot)$). It can be formulated as follows:
@@ IV Proposed Method: SparseInfer
Figure 2: Distribution of $X$, $W_{gate,i}$, and $Y=X\cdot W_{gate,i}$ values obtained from ProSparse-Llama2-13B during 8-shot inference on the GSM8K dataset.
@@ IV-A Training-free Sparsity Prediction
The reason why this approximation is relatively accurate is that it relies on the assumption that the elements of $X$ and $W_{gate}$ follow independent Gaussian distributions. The dot product of two vectors following independent Gaussian distributions will have a distribution close to that of a Laplace distribution, with a mean approximately equal to zero [20]. In this distribution, the ratio of positive to negative values is generally similar, meaning that the probability of individual products being positive or negative is almost equal. As a result, if there are more positive or negative values, the cumulative result is highly likely to have the sign of the more frequent values. This allows us to predict whether the result of $X\cdot W_{gate}$ will be positive or negative with relatively high accuracy, even without knowing the exact values.
Fig. 2 illustrates the distribution of values in a row vector of $W_{gate}$ ($W_{gate,i}$) and $X$ obtained from the ProSparse-Llama2-13B model [8] when performing 8-shot inference on a subset of the GSM8K dataset[21]. Due to space constraints, intermediate layers exhibiting similar distributions have been omitted. Both $X$ and $W_{gate}$ exhibit distributions that closely resemble Gaussian distributions, with an approximately equal ratio of positive and negative values. Moreover, their product $Y=X\cdot W_{gate,i}$ follows a Laplace distribution, as previously hypothesized. This confirms the validity of our initial assumption regarding the distributions of $X$ and $W_{gate}$, and consequently, the accuracy of our approximation method. In the early layers, the distribution of $X$ values is dominated by near-zero negative and near-zero positive values, leading to a narrow distribution heavily concentrated around zero, rather than following a typical normal distribution. This characteristic is believed to contribute to prediction inaccuracies.
@@
Fig. 3 shows the precision and recall of the proposed activation sparsity prediction for each layer in the ProSparse-Llama2-7B and ProSparse-Llama2-13B models. Here, precision represents the proportion of elements in $h1$ that the predictor classified as sparse and are indeed sparse (i.e., the likelihood that an element predicted to be sparse is actually sparse). On the other hand, On the other hand, recall indicates the proportion of truly sparse elements in $h1$ that the predictor successfully identified as sparse, relative to the total number of sparse elements (i.e., how well the predictor captures the overall sparsity). Overall, both models exhibited high precision values, reaching over 99% but a relatively lower precision was observed in the early layers compared to the later layers.
@@ V-A1 Latency
To compare the prediction overhead of SparseInfer with PowerInfer, the predictor latency with the first 10 samples in GSM8K dataset was measured on these inference engines. PowerInfer performs inference with PowerInfer-ProSparse-Llama2-13B model where the weights of the DEJAVU predictor trained with the rank size of 1024 is also included. Table I shows the number of operations in the PowerInfer predictor is even larger than the actual computation amount for MLP block. SparseInfer predictor has about an order of magnitude less number of operations than PowerInfer, and note that the operation is 32-bit bitwise XOR in SparseInfer while it is FP16 multiplication in PowerInfer. As a result, the predictor latency of SparseInfer for one token generation per layer is only 70 usec on average, which is 3.66$\times$ faster than that of PowerInfer. The reason why the speedup is less than the reduction ratio in the number of operations is that FP16 multiplications in PowerInfer are executed on Tensor cores while XORs are executed on CUDA cores.
@@ V-B End-to-End Speedup
With $\alpha$ being 1.00, SparseInfer achieved 1.79×1.79$\times$1.79 × and 1.74×1.74$\times$1.74 × speedups over llama.cpp for 13B and 7B models, respectively, outperforming PowerInfer by 1.27$\times$ and 1.30$\times$. As $\alpha$ gets larger, the speedups become slightly lower, due to the increased non-zero elements. With $\alpha$ being 1.02 for the 13B model, it leads to 1.70$\times$ over llama.cpp and 1.21$\times$ over PowerInfer. With $\alpha$ being 1.03 for the 7B model, it leads to 1.59$\times$ over llama.cpp and 1.19$\times$ over PowerInfer. We applied $\alpha>1.0$ only to the first 20 layers for both the 13B and 7B models.
The gain from the kernel fusion (+KF) turned out to be insignificant as the reused input vector size, which is 16KB and 20KB for 7B and 13B models respectively, is relatively small compared to the 4MB of L2 cache size per SM. Utilizing actual sparsity (+AS), in addition to the predicted activation sparsity, contributes significantly to the speedup, especially when $\alpha$ gets larger as it conservatively predicts the sparsity.
@@ V-C Accuracy
To evaluate the effect of the proposed approach on model performance, lm-harness[23], a widely used benchmark suite was executed. As SparseInfer utilizes activation sparsity only in the decoding phase, not in the prefill phase, log-likelihood-based benchmark evaluation is inadequate. We evaluated the model performance of ProSparse-Llama2-7B, 13B with GSM8K, BBH [24] benchmarks that generate the answer. Since a set of widely used benchmarks do not yet directly support llama.cpp, we ported the SparseInfer predictor and sparse GEMV to the PyTorch-based Transformer inference engine.

### v2

@@ Abstract
Leveraging sparsity is crucial for optimizing large language model (LLM) inference; however, modern LLMs employing SiLU as their activation function exhibit minimal activation sparsity. Recent research has proposed replacing SiLU with ReLU to induce significant activation sparsity and showed no downstream task accuracy degradation through fine-tuning. However, taking full advantage of it required training a predictor to estimate this sparsity. In this paper, we introduce SparseInfer, a simple, light-weight, and training-free predictor for activation sparsity of ReLU-fied LLMs, in which activation sparsity is predicted by comparing only the sign bits of inputs and weights. To compensate for possible prediction inaccuracy, an adaptive tuning of the predictor’s conservativeness is enabled, which can also serve as a control knob for optimizing LLM inference. The proposed method achieves approximately 21% faster inference speed over the state-of-the-art, with negligible accuracy loss of within 1%p.
@@ I Introduction
Following the remarkable success of Generative Pre-trained Transformer (GPT) models [1] based on the Transformer architecture [2], Large Language Models (LLMs) are being actively employed across a wide range of fields. The inference of LLMs based on the GPT architecture first performs an initial forward pass, also known as prefill, followed by an autoregressive decoding process that generates tokens sequentially, one at a time. Due to this sequential nature, memory accesses occur frequently in the decoding stage. This can lead to I/O bottlenecks, which in turn limits the overall computational performance. Well-known optimization methods such as pruning [3] or quantization [4] have been applied to LLMs to reduce this memory access overhead but can result in an accuracy loss.
@@ II Related Work
Activation Sparsity-induced LLM: Mirzadeh et al.[7] proposed a technique called ReLUfication which replaces activation functions in LLM, such as SiLU, with ReLU. They showed ReLUfication can induce significant activation sparsity while accuracy loss is negligible after fine-tuning. ProSparse [8] significantly induced activation sparsity by progressively applying sparsity regularization after Relufication. It also sets a positive value instead of zero for a ReLU threshold and achieves larger sparsity using a technique called FATReLU [11]. It induced a 22%p higher sparsity compared to ReluLLaMa [12], an open-source ReLU-based LLM fine-tuned from Llama2, achieving 66% sparsity. Despite a slight performance drop in some common sense benchmarks, the resultant model generally maintained comparable performance on average and even performed slightly better in certain cases, achieving up to a 4.65×4.65$\times$4.65 × speedup in inference time with PowerInfer [13]. A similar approach, TurboSparse [14], was proposed for Mixtral models.
@@ III Background: Gate-Based MLP
The gate-based MLP block consists of three linear layers ($W_{gate}$, $W_{up}$, and $W_{down}\in\mathbb{R}^{d\times k}$ in which $k>d$) with $d$ being the model dimension and one activation layer ($\sigma(\cdot)$). It can be formulated as follows:
@@ IV Proposed Method: SparseInfer
Figure 2: Distribution of $X$, $W_{gate,i}$, and $Y=X\odot W_{gate,i}$ values obtained from ProSparse-Llama2-13B during 8-shot inference on the GSM8K dataset.
@@ IV-A Training-free Sparsity Prediction
This approximation is relatively accurate because it relies on the assumption that the elements of $X$ and row vector of $W_{gate}$ ($W_{gate,i}$) follow independent Gaussian distributions with a mean equal to zero [15, 20]. It is known that the products of two variables following independent Gaussian distributions with zero means have a symmetric distribution with a mean equal to zero[21]. In this distribution, the ratio of positive to negative values is generally similar, meaning that the probability of individual products being positive or negative is almost equal. As a result, if there are more positive or negative values, the cumulative result is highly likely to have the sign of the more frequent values. This allows us to predict whether the result of $X\cdot W_{gate}$ will be positive or negative with relatively high accuracy, even without knowing the exact values.
Fig. 2 illustrates the distribution of values in a row vector of $W_{gate}$ ($W_{gate,i}$) and $X$ obtained from the ProSparse-Llama2-13B model [8] when performing 8-shot inference on a subset of the GSM8K dataset[22]. Due to space constraints, intermediate layers exhibiting similar distributions have been omitted.
Both $X$ and $W_{gate}$ exhibit distributions that closely resemble Gaussian distributions, with an approximately equal ratio of positive and negative values. Moreover, their products $Y=X\odot W_{gate,i}$ follow a symmetric distribution with a mean approaching zero, as previously hypothesized. This confirms the validity of our initial assumption regarding the distributions of $X$ and $W_{gate}$, and consequently, the accuracy of our approximation method. In the early layers, the distribution of $X$ values is dominated by near-zero negative and near-zero positive values, leading to a narrow distribution heavily concentrated around zero, rather than following a typical normal distribution. This characteristic is believed to contribute to prediction inaccuracies.
@@
Fig. 3 shows the precision and recall of the proposed activation sparsity prediction for each layer in the ProSparse-Llama2-7B and ProSparse-Llama2-13B models. Here, precision represents the proportion of elements in $h1$ that the predictor classified as sparse and are indeed sparse (i.e., the likelihood that an element predicted to be sparse is actually sparse). On the other hand, recall indicates the proportion of truly sparse elements in $h1$ that the predictor successfully identified as sparse, relative to the total number of sparse elements (i.e., how well the predictor captures the overall sparsity). Overall, both models exhibited high precision values, reaching over 99% but a relatively lower precision was observed in the early layers compared to the later layers.
@@ V-A1 Latency
To compare the prediction overhead of SparseInfer with PowerInfer, the predictor latency with the first 10 samples in GSM8K dataset was measured on these inference engines. PowerInfer performs inference with PowerInfer-ProSparse-Llama2-13B model where the weights of the DEJAVU predictor trained with the rank size of 1024 is also included. Table I shows the number of operations in the PowerInfer predictor is even larger than the actual computation amount for MLP block. SparseInfer predictor has about an order of magnitude less number of operations than PowerInfer, and note that the operation is 32-bit bitwise XOR in SparseInfer while it is FP16 multiplication in PowerInfer. As a result, the predictor latency of SparseInfer for one token generation per layer is only 70 usec on average, which is 3.66$\times$ faster than that of PowerInfer. The speedup is lower than the operation reduction ratio because FP16 multiplications in PowerInfer run on Tensor cores, while XORs run on CUDA cores.
@@ V-B End-to-End Speedup
With $\alpha$ being 1.00, SparseInfer achieved 1.79×1.79$\times$1.79 × and 1.74×1.74$\times$1.74 × speedups over llama.cpp for 13B and 7B models, respectively, outperforming PowerInfer by 1.27$\times$ and 1.30$\times$. As $\alpha$ gets larger, the speedups become slightly lower, due to the increased non-zero elements. With $\alpha$ being 1.03 for the 7B model, it leads to 1.59$\times$ over llama.cpp and 1.19$\times$ over PowerInfer. We applied $\alpha>1.0$ only to the first 20 layers for both the 13B and 7B models.
The gain from the kernel fusion (+KF) turned out to be insignificant due to the relatively small size of the input vector compared to the large cache size. Utilizing actual sparsity (+AS), in addition to the predicted activation sparsity, contributes significantly to the speedup, especially when $\alpha$ gets larger as it conservatively predicts the sparsity.
@@ V-C Accuracy
To evaluate the effect of the proposed approach on model performance, lm-harness[24], a widely used benchmark suite was executed. As SparseInfer utilizes activation sparsity only in the decoding phase, not in the prefill phase, log-likelihood-based benchmark evaluation is inadequate. We evaluated the model performance of ProSparse-Llama2-7B, 13B with GSM8K, BBH [25] benchmarks that generate the answer. Since a set of widely used benchmarks do not yet directly support llama.cpp, we ported the SparseInfer predictor and sparse GEMV to the PyTorch-based Transformer inference engine. Additionally, we were unable to measure accuracy because PowerInfer could not be ported to the Transformer engine.